OpenStudy (anonymous):

Can three vectors in the xy plane have u · v < 0 and v · w < 0 and u · w < 0?

OpenStudy (anonymous):

I don't think so

OpenStudy (anonymous):

It depends on your plane and what the vectors are measuring

OpenStudy (anonymous):

Dot product less than zero means obtuse angle

OpenStudy (anonymous):

all three angle can't be obtuse on one plane

OpenStudy (anonymous):

angle between u and v angle between u and w angle between w and v

OpenStudy (anonymous):

all are obtuse

OpenStudy (anonymous):

soooorrrryy i forgot we were talking about vectors hehe

OpenStudy (anonymous):

So since they would all be obtuse, it wouldn't be possible?

OpenStudy (anonymous):

However, for instance, Take: u = (1, 0), v = (−1/2, (square root of 3)/2), w = (−1/2, − (square root of 3)/2). Then u • v = v • w = u • w = −1/2 Which would, in fact, make it possible....right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Okay, I thought so.

OpenStudy (anonymous):

When figuring out this problem the key is to ask if three obtuse angle are possible on a plane. So if we arbitrarily choose two obtuse angel say 100 degree and 120 degree. Subtract the sum of the angel from 360 and we get an obtuse third angle.